OpenStudy (anonymous):

How do I solve this by completing the square: (r^2) - 2r - 1 = 0

OpenStudy (anonymous):

first thing move the -1 to the RHS add on both sides the square of half of 2 what do you get?

OpenStudy (anonymous):

(r^2) - 2r = 1

OpenStudy (anonymous):

That is all I understand so far

OpenStudy (anonymous):

what's half of 2?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

(r^2) - 2r + (1^2) = 1 + (1^2)

OpenStudy (anonymous):

good r^2-2r+1=2 the left handside is a remarkable identity (r-1)^2=2 can you now solve it?

OpenStudy (anonymous):

Do you just square the RHS?

OpenStudy (anonymous):

no i just added

OpenStudy (anonymous):

No, I mean is that what you do next?

OpenStudy (anonymous):



OpenStudy (anonymous):

i hope you know that if x^2=a then x=sqrt(a) or x=-sqrt(a)

OpenStudy (anonymous):

good,you got it

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

yw,don't forget to move the -1 to the RHS

OpenStudy (anonymous):

Yep :)